OpenStudy (anonymous):

Whats the value of N? 3n+6/3=15

OpenStudy (itsmichelle29):

n=13/3

OpenStudy (itsmichelle29):

welcome

OpenStudy (itsmichelle29):

3*n+6/3 = 15 // - 15 3*n+6/3-15 = 0 3*n-13 = 0 // + 13 3*n = 13 // : 3 n = 13/3 n = 13/3